# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

### v1

@@ Abstract
Large-scale pre-training using videos has proven effective for robot learning. However, the models pre-trained on such data can be suboptimal for robot learning due to the significant visual gap between human hands and those of different robots. To remedy this, we propose H2R, a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric videos. This process explicitly bridges the visual gap between human and robot embodiments during pre-training. We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2, replacing human hands with simulated robotic arms to generate robot-centric training data. Based on this, we construct and release a family of 1M-scale datasets covering multiple robot embodiments (UR5 with gripper/Leaphand, Franka) and data sources (SSv2, Ego4D). To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions. We validate H2R across three simulation benchmarks—Robomimic, RLBench, and PushT and real-world manipulation tasks with a UR5 robot equipped with Gripper and Leaphand end-effectors. H2R consistently improves downstream success rates, yielding gains of 5.0%–10.2% in simulation and 6.7%–23.3% in real-world tasks across various visual encoders and policy learning methods. These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains.
Figure 1: Overview of H2R. H2R is a data augmentation technique designed to enhance robot pre-training by converting first-person human hand operation videos into robot-centric visual data. By bridging the visual domain gap, H2R improves pre-trained visual encoders for downstream robot policies (imitation learning), validated across simulation benchmarks and real-world robotic tasks.
@@ 1 Introduction
Pre-training of generalizable robotic features for object manipulation and motion navigation constitutes a crucial objective within the realm of robotics. Inspired by the remarkable accomplishments of large scale pre-training in computer vision he2022mae ; nair2022r3m ; zeng2024learningmanipulationpredictinginteraction ; karamcheti2023voltron ; datacomp and natural language processing devlin2018bert ; yu2022pointbert ; openai2024gpt4technicalreport ; anil2023palm2technicalreport ; li2024datacomplm , many efforts have been devoted to harness large-scale data to construct generalizable representations in the robotics field embodimentcollaboration2024openxembodimentroboticlearning ; brohan2023rt1roboticstransformerrealworld ; brohan2023rt2visionlanguageactionmodelstransfer . Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive duan2023ar2 ; jang2022bczzeroshottaskgeneralization ; brohan2023rt1roboticstransformerrealworld ; khazatsky2024droidlargescaleinthewildrobot ; fang2023rh20tcomprehensiveroboticdataset ; gao2024efficientdatacollectionrobotic ; lee2021pickandplacetacklingroboticstacking ; herzog2023deeprlscalesorting ; kalashnikov2021mtoptcontinuousmultitaskrobotic ; duan2024manipulate ; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
@@
Recent works nair2022r3m ; xiao2022mvp ; zeng2024learning analyze such egocentric human video datasets such as Ego4D grauman2022ego4d , SSv2 goyal2017something , and Epic Kitchens damen2018scaling with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations. However, during the representation learning, the gap in visual representations between the human arm and the robotic arm remains largely unaddressed and can hinder the transferability of models trained on egocentric datasets to robotic systems. Specifically, when utilizing the robot expert data to fine-tune the pre-trained robotic representations for downstream robotic tasks, the model has to learn to bridge the visual gap between the first-person human hand and the robots in addition to acquiring nuanced task-specific skills demonstrated in the robot expert data. This would result in increased complexity during the fine-tuning process and suboptimal performance.
@@
To mitigate this issue, we propose H2R (as shown in Figure 1), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation. H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video. Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR pavlakos2023reconstructinghands3dtransformers to accurately detect the position and posture of the human hand in egocentric videos. Then, we simulate the same robot state in simulators to obtain the mask of robot actions. In the second stage, we use the Segment Anything Model kirillov2023segment to automatically separate human hand from background, and use the inpainting model LaMa suvorov2021resolutionrobustlargemaskinpainting to fill the removed hand mask. After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video. Finally, we overlay the robotic arm images captured by the simulator’s camera onto the areas where the human hands are removed. Through such a process, H2R explicitly reduces the gap between human and robot hands by creating realistic robotic arm movements that visually mimic human hand actions. It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
@@
Based on this pipeline, we construct and release a series of large-scale robot-centric datasets (H2R-1M), each containing approximately 1 million augmented frames. These datasets cover multiple robot embodiments (UR5 with gripper/Leaphand, Franka) and human video sources (SSv2, Ego4D), and provide modular annotations to support downstream robotic learning. To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics. This provides a lightweight and scalable proxy to assess the alignment quality between input human videos and robot-augmented outputs.
@@
To further verify the utility of the augmented datasets, we conduct downstream experiments comparing models pre-trained on original egocentric datasets with those trained on our released H2R-enhanced datasets. These comparisons are performed on both simulation and real-world robotic manipulation tasks to assess how H2R impacts policy learning performance in practice. For the pre-training stage, we apply MAE he2022mae and R3M nair2022r3m frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video. The pretrained visual representations are used in downstream imitation learning pipelines. We conduct comprehensive evaluations across both simulation and real-world manipulation tasks. In simulation, we evaluate pre-trained visual encoders on three benchmark suites: Robomimic, RLBench, and PushT. H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, and +5.3% for MAE and +7.0% for R3M on the PushT task. In real-world experiments, we deploy H2R-enhanced encoders on a UR5 robot equipped with both Gripper and Leaphand end-effectors, evaluated over six manipulation tasks under two policy frameworks: Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act . H2R brings consistent improvements, with average success rate gains of +13.3% (MAE) and +17.0% (R3M) under DP, and +10.8% (MAE) and +17.5% (R3M) under ACT. H2R also improves MAE/R3M performance when pre-trained on Ego4D, with up to +10% and +15% gains in simulated and real-world tasks, respectively. These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulated and real-world robotic manipulation settings.
Our paper provides three contributions:
  • We propose a data-centric pipeline, H2R, to mitigate the gap between human and robot hands when utilizing large-scale egocentric video datasets to pre-train generalizable visual features for robots.
  • We construct and release diverse robot-centric datasets (H2R-1M), combining multiple embodiments and egocentric video sources to support robot-compatible visual pretraining.
  • We demonstrate the effectiveness of H2R through extensive experiments on closed-loop benchmarks.
@@ 2 Related Work
Robot Imitation Learning. Data-driven policy learning sergey2018learning ; luo2024hilserl ; xiao2022mvp ; yang2024equibotsim3equivariantdiffusionpolicy ; chi2023diffusionpolicy ; embodimentcollaboration2024openxembodimentroboticlearning has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation. Imitation learning chi2023diffusionpolicy ; ze2024dp3 ; zhao2023act ; mandlekar2021matters trains policies from successful demonstrations, often supervised by behavior cloning torabi2018behavioral ; florence2021implicit objectives. ACT zhao2023act addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models chi2023diffusionpolicy ; ze2024dp3 are introduced to handle the inherent multimodality of robot motions. In addition, CordViP fu2025cordvip leverages 3D object-robot correspondences to enhance dexterous manipulation.
Visual Encoder Pretraining for Robotics. Visual pretraining improves generalization of robotic policies across diverse tasks. Researchers have explored architectural designs He_2016_CVPR ; dosovitskiy12021an , training objectives He_2022_CVPR ; chen2020simple ; chen2020big , and dataset compositions deng2009imagenet ; lin2014microsoft ; shao2019object365 ; Schuhmann2021LAION400MOD . PVR-Control parisi22a shows that pretrained visual representations can outperform direct state-based policies. RPT radosavovic2023robot tokenizes observations to enable masked prediction pretraining. Methods like MVP radosavovic2022realworld and R3M nair2022r3m utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning. Voltron karamcheti2023voltron demonstrates the use of MAE and contrastive learning for hierarchical robot control.
Cross-Domain Visual Alignment. Bridging the domain gap between human and robot visual inputs remains a major challenge. WHIRL bahl2022humantorobotimitationwild matches task structure from third-person views, while RoVi-Aug chen2024roviaugrobotviewpointaugmentation and Mirage chen2024mirage manipulate appearance via segmentation or image-space preprocessing. EgoMimic kareer2024egomimicscalingimitationlearning removes hands and normalizes views to align egocentric perspectives.
@@ 3 H2R: Human-to-Robot Data Augmentation
In this section, we describe H2R, a data augmentation pipeline for robot learning from egocentric human videos (Figure 2). It replaces human hands in every frame with robotic arms equipped with various end effectors, generating a new, visually different dataset. This approach aims to mitigate the visual gap between human hands and robots, thereby improving the generalizability of visual representations learned from egocentric data to robotic domains. Figure 3 illustrates examples of human hand videos that have been processed using H2R. These examples include different types of robots (UR5 and Franka) equipped with a variety of end-effectors, such as dexterous hands and grippers. We further introduce a CLIP-based metric to evaluate semantic consistency, and construct a family of large-scale datasets that combine different robot embodiments (UR5 with gripper, Leaphand, or both; Franka with gripper) and egocentric sources (SSv2 and Ego4D).
@@ 3.1 H2R Data Augmentation Pipeline
3D Hand Pose Estimation. In order to overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information. We adopt HaMeR pavlakos2023reconstructinghands3dtransformers , a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal. Given an ego-centric RGB image, HaMeR estimates both the hand pose (including 3D keypoints) and the intrinsic and extrinsic parameters of the rendering camera.
Human Arm and Hand Removal. After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background. We use Segment Anything Model (SAM) kirillov2023segment to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR. Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa suvorov2021resolutionrobustlargemaskinpainting , a powerful inpainting model, to fill in the removed hand-arm region. This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
robotic arm and End Effector Construction. This step involves constructing the robotic arm and different end effectors. There are two common types of end-effectors in robotic manipulation: grippers and dexterous hands. In our pipeline, we handle them differently based on their structural characteristics. For dexterous hands, which have multiple degrees of freedom and resemble human hands, we compute the joint angles by analyzing the hand keypoints predicted by HaMeR. Specifically, each finger joint angle is determined by the angle formed between three consecutive keypoints corresponding to that finger segment. For grippers, which typically consist of two or more parallel fingers, we determine the degree of opening or closing based on the Euclidean distance between the relevant fingertips in the hand keypoints. However, since hand keypoints alone are insufficient to fully define the robotic arm’s complete configuration, especially for joints that do not directly correspond to hand movement, we manually assign plausible values to these undetermined joint positions to construct a reasonable arm pose.
@@
A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix  B.
Robot Hand Rendering and Copy-paste. After setting the camera, the segmentation mask of the robotic arm is obtained by shooting with the camera. We directly obtain the pixel coordinates of the human hand keypoints from HaMeR, which are predicted from the input image. In parallel, the pixel coordinates of the robot end-effector links are computed in the simulator by projecting their 3D positions through the aligned camera using the known transformation matrices. By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.

### v3

@@ Abstract
Large-scale pre-training using videos has proven effective for robot learning. However, the models pre-trained on such data can be suboptimal for robot learning due to the significant visual gap between human hands and those of different robots. To remedy this, we propose H2R, a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric videos. This process explicitly bridges the visual gap between human and robot embodiments during pre-training. We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2, replacing human hands with simulated robotic arms to generate robot-centric training data. To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions. We validate H2R across four simulation benchmarks—Robomimic, RLBench, PushT and CortexBench, and real-world manipulation tasks with UR5 and Dual-Arm Franka/UR5. H2R consistently improves downstream success rates, yielding gains of 1.3%–10.2% in simulation and 3.3%–23.3% in real-world tasks across various visual encoders and policy learning methods. We further demonstrate the potential of H2R in cross-embodiment generalization and its compatibility with VLA models. These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains. The project page is available in https://sites.google.com/view/h2r-robotics.
@@ 1 Introduction
Pre-training of generalizable robotic features for object manipulation and motion navigation constitutes a crucial objective within the realm of robotics. Inspired by the remarkable accomplishments of large scale pre-training in computer vision [he2022mae, nair2022r3m, zeng2024learningmanipulationpredictinginteraction, karamcheti2023voltron, datacomp] and natural language processing [devlin2018bert, yu2022pointbert, openai2024gpt4technicalreport, anil2023palm2technicalreport, li2024datacomplm], many efforts have been devoted to harnessing large-scale data to construct generalizable representations in the robotics field [embodimentcollaboration2024openxembodimentroboticlearning, brohan2023rt1roboticstransformerrealworld, brohan2023rt2visionlanguageactionmodelstransfer]. Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive [duan2023ar2, jang2022bczzeroshottaskgeneralization, brohan2023rt1roboticstransformerrealworld, khazatsky2024droidlargescaleinthewildrobot, fang2023rh20tcomprehensiveroboticdataset, gao2024efficientdatacollectionrobotic, lee2021pickandplacetacklingroboticstacking, herzog2023deeprlscalesorting, kalashnikov2021mtoptcontinuousmultitaskrobotic, duan2024manipulate]; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
@@
Recent works [nair2022r3m, xiao2022mvp, zeng2024learning] analyze such egocentric human video datasets such as Ego4D [grauman2022ego4d], SSv2 [goyal2017something], and Epic Kitchens [damen2018scaling] with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations. However, during the representation learning, the gap in visual representations between the human arm and the robotic arm remains largely unaddressed and can hinder the transferability of models trained on egocentric datasets to robotic systems. Specifically, when utilizing the robot expert data to fine-tune the pre-trained robotic representations for downstream robotic tasks, the model has to learn to bridge the visual gap between the first-person human hand and the robots in addition to acquiring nuanced task-specific skills demonstrated in the robot expert data. This would result in increased complexity during the fine-tuning process and suboptimal performance.
@@
To mitigate this issue, we propose H2R (as shown in Figure LABEL:fig:pipeline), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation. H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video. Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR [pavlakos2023reconstructinghands3dtransformers] to accurately detect the position and posture of the human hand in egocentric videos. Then, we simulate the same robot state in simulators to obtain the mask of robot actions. In the second stage, we use the Segment Anything Model [kirillov2023segment] to automatically separate human hand from background, and use the inpainting model LaMa [suvorov2021resolutionrobustlargemaskinpainting] to fill the removed hand mask. After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video. Finally, we overlay the robotic arm images captured by the simulator’s camera onto the areas where the human hands are removed. Through such a process, H2R explicitly reduces the gap between human and robot hands by creating realistic robotic arm movements that visually mimic human hand actions. It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
@@
To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics. This provides a lightweight and scalable proxy to assess the alignment quality between input human videos and robot-augmented outputs.
@@
To further verify the utility of the augmented datasets, we conduct downstream experiments comparing models pre-trained on original egocentric datasets with those trained on our released H2R-enhanced datasets. These comparisons are performed on both simulation and real-world robotic manipulation tasks to assess how H2R impacts policy learning performance in practice. For the pre-training stage, we apply MAE [he2022mae] and R3M [nair2022r3m] frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video. The pretrained visual representations are used in downstream imitation learning pipelines. We conduct comprehensive evaluations across both simulation and real-world manipulation tasks. In simulation, we evaluate pre-trained visual encoders on four benchmark suites: Robomimic, RLBench, PushT and CortexBench. H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, +5.3% for MAE and +7.0% for R3M on the PushT task, and +5.3% for MAE and +1.3% for R3M on the CortexBench task. In real-world experiments, we deploy H2R-enhanced encoders on different robot and end-effector settings, evaluated over nine manipulation tasks under two policy frameworks: Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act]. H2R brings consistent improvements, with average success rate gains of +10.0% (MAE) and +14.1% (R3M) under DP, and +12.8% (MAE) and +13.3% (R3M) under ACT. H2R also improves MAE/R3M performance when pre-trained on Ego4D, with up to +10% and +15% gains in simulated and real-world tasks, respectively. Beyond these evaluations, we also test H2R’s ability to support cross-embodiment transfer by pre-training encoders on Franka data and deploying them on UR robots, indicating that H2R improves representation transfer across different morphologies. In addition, fine-tuning VLA backbones on H2R-generated datasets yields markedly higher success rates on dual-arm UR tasks than fine-tuning on human-hand videos or using no fine-tuning. These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulation and real-world manipulation settings.
@@ 2 Related Work
Robot Imitation Learning. Data-driven policy learning [sergey2018learning, luo2024hilserl, xiao2022mvp, yang2024equibotsim3equivariantdiffusionpolicy, chi2023diffusionpolicy, embodimentcollaboration2024openxembodimentroboticlearning] has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation. Imitation learning [chi2023diffusionpolicy, ze2024dp3, zhao2023act, mandlekar2021matters] trains policies from successful demonstrations, often supervised by behavior cloning [torabi2018behavioral, florence2021implicit] objectives. ACT [zhao2023act] addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models [chi2023diffusionpolicy, ze2024dp3] are introduced to handle the inherent multimodality of robot motions. In addition, CordViP [fu2025cordvip] leverages 3D object-robot correspondences to enhance dexterous manipulation.
Visual Encoder Pretraining for Robotics. Visual pretraining improves generalization of robotic policies across diverse tasks. Researchers have explored architectural designs [He_2016_CVPR, dosovitskiy12021an], training objectives [He_2022_CVPR, chen2020simple, chen2020big], and dataset compositions [deng2009imagenet, lin2014microsoft, shao2019object365, Schuhmann2021LAION400MOD]. PVR-Control [parisi22a] shows that pretrained visual representations can outperform direct state-based policies. RPT [radosavovic2023robot] tokenizes observations to enable masked prediction pretraining. Methods like MVP [radosavovic2022realworld] and R3M [nair2022r3m] utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning. Voltron [karamcheti2023voltron] demonstrates the use of MAE and contrastive learning for hierarchical robot control.
Cross-Domain Visual Alignment. Bridging the domain gap between human and robot visual inputs remains a major challenge. WHIRL [bahl2022humantorobotimitationwild] matches task structure from third-person views, while RoVi-Aug [chen2024roviaugrobotviewpointaugmentation] and Mirage [chen2024mirage] manipulate appearance via segmentation or image-space preprocessing. EgoMimic [kareer2024egomimicscalingimitationlearning] removes hands and normalizes views to align egocentric perspectives.
@@ 3 H2R: Human-to-Robot Data Augmentation
In this section, we describe H2R, a data augmentation pipeline for robot learning from egocentric human videos (Figure 1). It replaces human hands in every frame with robotic arms equipped with various end effectors, generating a new, visually different dataset. This approach aims to mitigate the visual gap between human hands and robots, thereby improving the generalizability of visual representations learned from egocentric data to robotic domains. We first examine the position and pose of the hand and arm in the image, then map this pose to the robotic arm and end-effector in the simulator. Next, we erase the original hand and arm from the image and overlay a rendered image of the simulated robot performing the same pose onto the original image in place of the removed hand and arm. Figure 3 illustrates examples of human hand videos that have been processed using H2R.
@@ 3.1 H2R Data Augmentation Pipeline
3D Hand Pose Estimation. In order to overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information. We adopt HaMeR [pavlakos2023reconstructinghands3dtransformers], a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal. Given an ego-centric RGB image, HaMeR estimates both the hand pose (including 3D keypoints) and the intrinsic and extrinsic parameters of the rendering camera.
Robotic Arm and End Effector Construction. This step involves building the robotic arm and its end-effectors, which are typically either grippers or dexterous hands. For dexterous hands, we estimate joint angles from hand keypoints predicted by HaMeR. Each finger joint angle is calculated using three consecutive keypoints along that finger. For grippers, we determine how open or closed they are based on the Euclidean distance between the corresponding fingertips. Since hand keypoints don’t capture the full arm pose (especially for joints unrelated to the hand), we manually set reasonable values for the remaining arm joints to complete a plausible robot configuration.
@@
A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix.
Robot Hand Rendering and Copy-paste. After setting the camera, the segmentation mask of the robotic arm is obtained by rendering with the simulator camera. We directly obtain the pixel coordinates of the human hand keypoints from HaMeR, which are predicted from the input image. In parallel, the pixel coordinates of the robot end-effector links are computed in the simulator by projecting their 3D positions through the aligned camera using the known transformation matrices. By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.
Human Arm and Hand Removal. After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background. We use Segment Anything Model (SAM) [kirillov2023segment] to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR. Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa [suvorov2021resolutionrobustlargemaskinpainting], a powerful inpainting model, to fill in the removed hand-arm region. This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.